Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 353     Entry time: 22 Sep 2023, 07:27
Author: Raphael de Wijn 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 7 keV, 202 pulses, 564 kHz, ~1.3 mJ, stable delivery

Optical laser delivery

  • None

Achievements/Observations

  • Collected a few NQO1 runs (no mixing) in Adaptive and FixedMG with the Ros setup, good diffraction
  • Difficult to reach droplet synchronisation
ELOG V3.1.4-7c3fd00